### v2

# SAEmnesia: Erasing Concepts in Diffusion Models with Supervised Sparse Autoencoders

Abstract
Concept unlearning in diffusion models is hampered by feature splitting, where concepts are distributed across many latent features, making their removal challenging and computationally expensive. We introduce SAEmnesia, a supervised sparse autoencoder framework that overcomes this by enforcing one-to-one concept-neuron mappings. By systematically labeling concepts during training, our method achieves feature centralization, binding each concept to a single, interpretable neuron. This enables highly targeted and efficient concept erasure. SAEmnesia reduces hyperparameter search by 96.7% and achieves a 9.2% improvement over the state-of-the-art on the UnlearnCanvas benchmark. Our method also demonstrates superior scalability in sequential unlearning, improving accuracy by 28.4% when removing nine objects, establishing a new standard for precise and controllable concept erasure. Moreover, SAEmnesia mitigates the possibility of generating unwanted content under adversarial attack and effectively removes nudity when evaluated with I2P.
1 Introduction
Text-to-image diffusion models have achieved remarkable success in generating high-quality images from textual descriptions, with applications across diverse domains [rombach2022high]. However, they can also produce harmful, inappropriate, or copyrighted content when given specific prompts, raising safety concerns. This has spurred growing interest in machine unlearning, which aims to selectively remove undesired concepts from trained models while preserving their generative abilities [zhang2024unlearncanvas].
Figure 1: SAEmnesia enables precise concept-level manipulation: each latent activates for a single concept (monosemanticity), and each concept is embedded in a single latent (feature centralization). So, to erase a target concept, we only need to steer a single latent. The removed concepts correctly disappear in the diagonal images (“Architectures”, “Bears”, “Sandwiches”) while the corresponding style is preserved. Note that they remain present in the non-diagonal ones, thereby preserving the fidelity and diversity when unlearning unrelated content.
A core challenge in concept unlearning is identifying where and how concepts are represented inside these models. Each neuron can encode multiple unrelated concepts simultaneously. This phenomenon is known as polysemanticity, making interpretability even more challenging. Mechanistic interpretability (MI) seeks to understand the internal workings of neural networks by analyzing their representations. Sparse Autoencoders (SAEs) provide a particularly effective MI tool by decomposing model activations into sparse and interpretable concept-level features [bricken2023monosemanticity]. In this work, we adopt the notion of features as the fundamental units of neural network representations that cannot be further decomposed into simpler independent factors, as defined by  bereska2024mechanistic. Neural networks can capture natural abstractions [Chan2023NaturalAbstractions] through their learned features, which serve as building blocks of their internal representations, aiming to capture the concepts underlying the data. For simplicity, we use the terms “concepts” and “features” interchangeably, as well as “neurons” and “latents”. SAEs aim to learn monosemantic latents, meaning that they activate almost exclusively for a specific concept. On the other hand, to improve models’ interpretability even further, individual concepts should not be spread across many latents, aiming for a one-to-one mapping. Yet, in practice, multiple latents often respond to the same concept, this phenomenon is also known as feature splitting [bricken2023monosemanticity]. This means modifying one concept requires changing multiple neurons. This distributed nature poses two key challenges: (i) exhaustive searches across many latent combinations to find the right subset to modify, making unlearning computationally expensive; and (ii) overlapping latents blur concept boundaries, so interventions risk unintended side effects on related concepts.
To overcome feature splitting, we introduce SAEmnesia, that enriches the SAE training framework enforcing one-to-one mappings between concepts and latents through supervised labeling. Therefore, our method achieves feature centralization, localizing each concept into a single latent and thereby preventing splitting across multiple neurons. This binding simplifies mechanistic unlearning, thanks to a precise single-latent intervention. Fig. 1 illustrates the effect of SAEmnesia for concept removal in generative models, demonstrating the possibility of precise control over what the model generates, while retaining overall quality and diversity. On the UnlearnCanvas benchmark [zhang2024unlearncanvas], SAEmnesia achieves an 9.22% improvement over the state-of-the-art mechanistic approaches [cywinski2025saeuron]. In sequential unlearning tasks, we demonstrate superior scalability with a 28.4% improvement in unlearning accuracy for 9-object removal. Furthermore, at inference, this interpretable representation reduces hyperparameter search by 96.67%.
In summary, our key contributions are as follows: (i) we introduce a supervised sparse autoencoder that explicitly enforces one-to-one concept–latent mappings, eliminating feature splitting and enabling transparent, interpretable control over concept representations; (ii) we show that this structure makes concept erasure significantly more efficient: each concept can be removed by steering a single latent, substantially reducing inference-time hyperparameter search; (iii) we achieve state-of-the-art performance on the UnlearnCanvas benchmark and demonstrate improved sequential unlearning, stronger robustness to adversarial attacks, and more effective NSFW-content suppression.
2 Related Work
Machine unlearning in diffusion models. Machine unlearning was first introduced by cao2015towards, who transformed neural network models through additional simple layers into formats where output is a summation of independent features, allowing unlearning by blocking selected summation weights or nodes. However, recent works focusing on unlearning for diffusion models typically employ fine-tuning approaches to unlearn specific concepts. EDiff [wu2024erasediff] formulates this problem as bi-level optimization, while ESD [gandikota2023erasing] leverages negative classifier-free guidance for concept removal. FMN [zhang2024forget] introduces a re-steering loss applied only to attention layers, and SalUn [fan2023salun] and SHS [wu2024scissorhands] select parameters to adapt through saliency maps or connection sensitivity. SA [heng2023selective] replaces unwanted data distribution with surrogate distributions, with an extension to selected anchor concepts in CA [kumari2023ablating]. SPM [lyu2024one] takes a different approach, using small linear adapters added after each linear and convolutional layer to directly block unwanted content propagation. Methods that do not rely on fine-tuning include SEOT [li2024get], which removes unwanted content from text embeddings, and UCE [gandikota2024unified], which adapts cross-attention weights using closed-form solutions. In contrast to these approaches, our work leverages SAEs to achieve unlearning through interpretable feature manipulation during inference, without modifying the base model weights and providing full transparency into which specific features are being targeted for removal.
Beyond individual methods, recent surveys synthesize objectives, taxonomies, and evaluation protocols for generative model unlearning, offering broader context for method design and assessment [feng2025surveygenmu, chen2025unlearning_survey]. Furthermore, recent analyses highlight instability and concept resurgence after unlearning (such as revival under subsequent fine-tuning or adversarial prompting) reinforcing the need for interpretable and stable interventions [george2025illusion, lu2025when, suriyakumar2024resurgence].
SAEs background. Our aim is to enable effective concept unlearning in diffusion models by selectively removing unwanted concepts while preserving generative quality. To achieve this, we decompose the high-dimensional, entangled activations from Stable Diffusion (SD) into sparse, interpretable feature directions that correspond to meaningful visual concepts. SAEs serve as the key tool for this decomposition, enabling us to map individual neurons to specific semantic concepts and subsequently intervene on them for targeted unlearning. A standard single-layer ReLU SAE [Olshausen1997SparseCW] operates on $d$-dimensional activation vectors. Let $\mathbf{x}\in\mathbb{R}^{d}$ denote the input activation vector and $n$ be the latent dimension, typically set to $d$ multiplied by a positive expansion factor. The encoder and decoder are defined as [bricken2023monosemanticity]:
$\begin{split}\mathbf{v}&=\text{ReLU}(\mathbf{W}_{\text{enc}}(\mathbf{x}-\mathbf{b}_{\text{pre}})+\mathbf{b}_{\text{enc}})\\ \hat{\mathbf{x}}&=\mathbf{W}_{\text{dec}}\mathbf{v}+\mathbf{b}_{\text{pre}},\end{split}$ (1)
where $\mathbf{v}$ is the sparse hidden representation, $\hat{\mathbf{x}}$ is the reconstructed input, $\mathbf{W}_{\text{enc}}\in\mathbb{R}^{n\times d}$ and $\mathbf{W}_{\text{dec}}\in\mathbb{R}^{d\times n}$ are the encoder and decoder weight matrices respectively, and $\mathbf{b}_{\text{pre}}\in\mathbb{R}^{d}$ and $\mathbf{b}_{\text{enc}}\in\mathbb{R}^{n}$ are learnable bias terms.
TopK SAEs. In our work, we employ TopK SAEs [Makhzani2013kSparseA] that provide enhanced sparsity control. The TopK activation function identifies the $k$ largest pre-activations and sets all others to zero, ensuring sparsity while preserving the most significant features:
$\begin{split}\mathbf{z}&=\text{TopK}(\mathbf{v}),\\ \hat{\mathbf{x}}&=\mathbf{W}_{\text{dec}}\mathbf{z}+\mathbf{b}_{\text{pre}}.\end{split}$ (2)
Here, $\mathbf{v}$ represents the pre-TopK activations, while $\mathbf{z}$ represents the post-TopK sparse activations.
Training objective. Given a mini-batch of size $B$, the TopK SAE loss combines reconstruction error with an auxiliary loss to prevent dead latents:
$\mathcal{L}_{\text{unsupSAE}}=\frac{1}{B}\sum_{b=1}^{B}\big\|\mathbf{x}^{(b)}-\hat{\mathbf{x}}^{(b)}\big\|_{2}^{2}+\alpha\,\mathcal{L}_{\text{aux}},$ (3)
where $\mathbf{x}-\hat{\mathbf{x}}||_{2}^{2}$ is the reconstruction error and $\mathcal{L}_{\text{aux}}$ is an auxiliary loss using only the largest $k_{\text{aux}}$ feature activations that have not fired on a large number of training samples (so-called dead latents). The auxiliary loss prevents dead latents from occurring and is scaled by coefficient $\alpha$.
Interpretability of vision models. SAEs have recently gained traction as a tool for uncovering human-interpretable structure in high-dimensional representations. Early applications focused on discriminative models, such as interpreting features in CLIP [fry2024towards, daujotas2024case, rao2024discover] or traditional classifiers [szegedy2015going, gorton2024missing]. More recent work has extended SAEs to vision–language settings, enabling tasks such as hallucination mitigation [jiang2024interpreting] and interpretable report generation [abdulaal2024x]. Across these domains, SAEs provide explicit concept–neuron mappings that move beyond post-hoc explanations and allow more precise control over learned representations. Unlike previous work on discriminative models, our approach uses SAEs with diffusion models to directly identify and control specific concepts.
Within diffusion models, interpretability research has primarily examined how semantic information propagates through the architecture. Studies have identified meaningful directions in UNet bottlenecks [kwon2022diffusion, park2023understanding, hahm2024isometric, ijishakin2024h, Kim_2025_ICCV], analyzed cross-attention to link prompts with spatial activations [tang2022daam, basu2023localizing, basu2024mechanistic], and even used intermediate text encoder states for generation [toker2024diffusion]. Although these advances shed light on the internals of the model, they generally do not enable targeted interventions. Our approach yields sparse representations that support both interpretation and controllable unlearning of specific concepts.
SAE-Based unlearning. Recent work has explored applying SAEs to diffusion models for concept manipulation. kim2025concept introduced Concept Steerers, training SAEs on text embedding representations to identify concept-specific directions before cross-attention processing. While their approach achieves effective concept manipulation, working only on text encoders can lead to suboptimal results, especially when facing adversarial attacks that exploit deeper model representations. cywinski2025saeuron introduced SAeUron, a post-cross-attention approach using SAEs trained on diffusion model activations in an unsupervised manner. While SAeUron achieves state-of-the-art performance on UnlearnCanvas, its unsupervised training creates weak concept-latent associations. This means that concepts are still represented across multiple neurons. This distributed representation requires computationally expensive feature threshold searches to identify which combination of latent features must be modified to unlearn each concept. In contrast, we only need to steer a single feature per concept.
3 Methodology
Figure 2: SAEmnesia pipeline. Training comprises two phases: (i) establishing sparse representations via standard unsupervised SAE training, (ii) applying supervised losses to strengthen specific concept-neuron associations. During inference, we need to steer a single latent per concept.
Unlike unsupervised SAE training methods [bricken2023monosemanticity] that require post-hoc discovery of concept-relevant features, our supervised approach directly enforces concept-latent assignments during training to achieve stronger one-to-one mappings (see Fig. 2). Although our approach requires supervision, the labels come at no additional cost, as they are directly derived from the text prompts used to generate or condition the images: i.e., the same concepts that the SAE aims to forget.
SAEmnesia for diffusion models. We apply SAEmnesia to diffusion models by training on activations extracted from every timestep $t$ of the denoising diffusion process. These activations are obtained from the cross-attention blocks of the diffusion model and form feature maps. Each feature map extracted at timestep $t$ is a spatially structured tensor of shape $\mathbf{F}_{t}\in\mathbb{R}^{h\times w\times d}$, where $h$ and $w$ denote the height and width of the feature map, and $d$ is the dimensionality of each feature vector. Each spatial position within the feature map corresponds to a patch in the input image. As a single SAE training sample, we consider an individual $d$-dimensional feature vector $\mathbf{x}$, disregarding the information about its spatial position. Therefore, from each feature map, we obtain $h\times w$ training samples. Note that our method is architecture agnostic, meaning that it can be transferred to various text-to-image (T2I) models.
Concept-latent assignment. Here we introduce supervised training that directly assigns concepts to specific latents. To determine which latent should be assigned to each concept during the supervised phase and to validate the quality of these assignments after training, we utilize the score function [cywinski2025saeuron], defined in Eq. 4, measuring feature-concept correspondence. Given a dataset of activations $D=D_{c}\cup D_{\neg c}$, where $D_{c}$ contains data of the target concept $c$ and $D_{\neg c}$ does not, the score function is defined as:
$\begin{split}\text{score}(i,t,c,D)&=\frac{\mu(i,t,D_{c})}{\sum_{j=1}^{n}\mu(j,t,D_{c})+\delta}\\ &\quad-\frac{\mu(i,t,D_{\neg c})}{\sum_{j=1}^{n}\mu(j,t,D_{\neg c})+\delta}\end{split}$ (4)
where $\delta$ prevents division by zero and $\mu(i,t,D)=\frac{1}{|D|}\sum_{x\in D}z_{i}$ denotes the average activation of the $i$-th feature on activations from timestep $t$ (we omit $t$ from $z_{i}$ for simplicity). Features with high scores exhibit strong activation for concept $c$ while remaining weakly activated for other concepts. Formally, in order to achieve feature centralization, for a given concept $c$, the score function score(i,t,c,D)(i,t,c,D) achieves a high value for a single latent index $i$, while remaining low for all other indices $j\neq i$:
$\text{score}(i,t,c,D)\gg\text{score}(j,t,c,D),\quad\forall j\neq i.$ (5)
Given a set $\mathcal{C}=\{c_{1},\ldots,c_{K}\}$ of concepts to unlearn, we define a mapping $\Phi:\mathcal{C}\to\{1,\ldots,n\}$ assigning each concept $c$ to the latent index $\Phi(c)=i_{c}$ with the highest score. For training samples containing multiple concepts, we assign multiple target indices corresponding to each present concept. By enforcing one-to-one concept-to-latent mappings, the cardinality of the SAE’s hidden layer directly corresponds to the number of concepts that can be explicitly represented and manipulated.
SAEmnesia loss function. We employ a composite loss function designed to maintain SAE reconstruction capabilities while strengthening concept-latent associations:
$\mathcal{L}_{\text{{SAEmnesia}}}=\mathcal{L}_{\text{unsupSAE}}+\beta\mathcal{L}_{\text{supSAE}}+\lambda\mathcal{L}_{\text{L1}}.$ (6)
Where $\mathcal{L}_{\text{unsupSAE}}$ is the loss as defined in Eq. 3, $\mathcal{L}_{\text{supSAE}}$ is the supervised loss that enforces the desired concept-latent relationships and $\mathcal{L}_{\text{L1}}$ is the sparsity regularization term. $\mathcal{L}_{\text{supSAE}}$ consist of our Concept Assignment (CA) loss with a weighted additional Decorrelation (DC) loss:
$\mathcal{L}_{\text{supSAE}}=\mathcal{L}_{\text{CA}}+\gamma\mathcal{L}_{\text{DC}}.$ (7)
Concept assignment loss. Unlike traditional approaches that only apply reconstruction loss globally across all latents, our method strengthen concept-latent bonds that enable one-to-one mappings. Our CA loss is computed exclusively for latents that are assigned to concepts present in each training sample, encouraging these specific latents to activate strongly. Given a training sample, we define the binary ground-truth vector $\mathbf{y}=[y_{1},\ldots,y_{K}]^{\top}\in\{0,1\}^{K}$, where $y_{k}=1$ if concept $c_{k}$ is present in the sample, and 0 otherwise. We denote with $\mathcal{T}$ the set of concepts present in that sample. The CA loss measures how much the assigned latents activate when their corresponding concepts are present:
$\mathcal{L}_{\text{CA}}=\frac{1}{B}\sum_{b=1}^{B}\frac{1}{|\mathcal{T}^{(b)}|}\sum_{c\in\mathcal{T}^{(b)}}\big[-\log\sigma(v^{(b)}_{i_{c}})\big$], (8)
where $v_{i_{c}}$ represents the pre-activation (logit), and $\sigma(\cdot)$ is the sigmoid function.
Decorrelation constraint. To promote disentanglement across multiple macro-categories of concepts, we generalize the decorrelation constraint to $M$ disjoint concept groups. Specifically, we partition the full concept set into non-overlapping subsets $\mathcal{C}=\bigcup_{m=1}^{M}\mathcal{C}_{m}$, with $\mathcal{C}_{m}\cap\mathcal{C}_{m^{\prime}}=\varnothing$ for $m\neq m^{\prime}$, where each $\mathcal{C}_{m}$ represents a high-level group of related concepts (e.g., objects, styles, materials, or other semantic categories depending on the dataset). We denote by $\mathcal{I}_{\mathcal{C}_{m}}=\{\Phi(c)\mid c\in\mathcal{C}_{m}\}$ the set of latent indices assigned to the concepts in group $\mathcal{C}_{m}$. Given a mini-batch of size $B$, we define the activation vector of each concept $c$ as $\mathbf{a}_{c}=[v^{(1)}_{i_{c}},v^{(2)}_{i_{c}},\ldots,v^{(B)}_{i_{c}}$$^{\top}$. The multi-group decorrelation constraint is then formulated as:
$\mathcal{L}_{\text{DC}}=\frac{\displaystyle\sum_{m<m^{\prime}}\sum_{i\in\mathcal{I}_{\mathcal{C}_{m}}}\sum_{j\in\mathcal{I}_{\mathcal{C}_{m^{\prime}}}}\rho(\mathbf{a}_{i},\mathbf{a}_{j})}{\displaystyle\sum_{m<m^{\prime}}|\mathcal{I}_{\mathcal{C}_{m}}|\,|\mathcal{I}_{\mathcal{C}_{m^{\prime}}}|}.$ (9)
Here, $\rho(\mathbf{a}_{i},\mathbf{a}_{j})$ denotes the Pearson correlation coefficient between activation vectors $\mathbf{a}_{i}$ and $\mathbf{a}_{j}$. This constraint penalizes correlations between activation patterns of latents assigned to different concept groups.
Sparsity regularization. To encourage sparse activations in the latent representation, we incorporate an L1 regularization term that penalizes the magnitude of latent activations $\mathbf{v}$. This sparsity constraint promotes the emergence of interpretable features by encouraging most latents to remain inactive for any given input, thereby improving the disentanglement of learned representations.
Feature centralization. As defined in Eq. 5, feature centralization occurs when only one feature achieves a high score for a specific concept. However, computing the score function requires the entire dataset, making it impractical during training. Instead, we use $\mathcal{L}_{\text{supSAE}}$ as a proxy to achieve feature centralization. $\mathcal{L}_{\text{CA}}$ directly enforces that assigned latents activate strongly for their target concepts, while $\mathcal{L}_{\text{L1}}$ maintains sparsity across all latents. This combination concentrates a concept’s information in the designated latent.
Inference and concept unlearning. To unlearn a concept $c$, SAEmnesia only needs the single latent $i_{c}$ to erase the concept. The activation of the selected feature is multiplied by a negative value $\gamma_{c}<0$ normalized by the average activation $\mu(i_{c},t,D_{c})$ on concept samples of a validation dataset $D$. This removes the influence of the targeted concept on the activation vector $\mathbf{z}$. Each $i_{c}$-th latent feature activation is modified as follows:
$z_{i_{c}}=\begin{cases}\gamma_{c}\mu(i_{c},t,D_{c})z_{i_{c}},&\text{if }z_{i_{c}}>\mu(i_{c},t,D)\\ z_{i_{c}},&\text{otherwise}\end{cases}$ (10)
The condition $z_{i_{c}}>\mu(i_{c},t,D)$ prevents random feature ablation when scores are low. During inference, we can use the original pretrained model for the first $t$ steps, setting the multipliers to 1 and retaining the pretrained model priors. We can then turn on SAEmnesia for the remaining steps.
4 Experiments and Results
We conduct comprehensive experiments to evaluate SAEmnesia across multiple dimensions: unlearning effectiveness, generation quality, concept separation, performance robustness, incremental unlearning capabilities, adversarial resilience and nudity removal. Note that our focus in this work is on object erasure, as it presents a considerably greater challenge than style erasure [cywinski2025saeuron].
4.1 Experimental setup
We report here the evaluation setup for the experiments of our proposed method, including datasets, architectures and evaluation metrics.
Dataset. We extract activations from SD v1.5 [rombach2022high] within the UnlearnCanvas Benchmark [zhang2024unlearncanvas]. SD v1.x family remains the standard evaluation setting for concept-erasure research: the benchmark itself is built on SD v1.5, and most contemporary methods report results on SD v1.x [cywinski2025saeuron, Thakral_2025_CVPR, Wang_2025_CVPR, ACE_2025_CVPR, Lee_2025_CVPR, Srivatsan_2025_CVPR]. This makes SD v1.x the appropriate and widely-accepted testbed for fair comparison and reproducibility. We construct labeled training data by generating activations from structured prompts with known concept compositions. For each object class $c$ (e.g., “Bears”, “Cats”) and style $s$ (e.g., “Impressionism”, “Cubism”), we generate 80 prompts of the form An image of {object} in {style} and collect feature maps from selected U-Net cross-attention blocks across all 50 denoising timesteps during text-conditioned generation. Each activation is directly labeled with its corresponding object and style based on the prompt structure, creating explicit concept-activation pairs. The employed objects and styles are taken from the UnlearnCanvas benchmark, which consists of 20 different objects and 50 different styles. We focus our analysis on block up.1.1 for object-related features, as this block has been empirically demonstrated to specialize in generating specific visual aspects [basu2023localizing]. This controlled labeling strategy provides clean supervision signals that enable direct concept-neuron mapping during SAE training, contrasting with unsupervised approaches that must discover concept representations through post-hoc analysis. For evaluation, we employed the same setting of SAeUron [cywinski2025saeuron].
SAE model. Our best-performing method starts with a pre-trained unsupervised SAE and fine-tunes it using SAEmnesia loss in Eq. 6. Unless otherwise stated, SAEmnesia is applied across all denoising steps for a fair comparison with [cywinski2025saeuron]. For the decorrelation constraint in Eq. 9, we choose $\mathcal{C}_{\text{obj}}$ and $\mathcal{C}_{\text{sty}}$, corresponding to object and style concepts, to enforce concept separation. Loss function hyperparameters and additional model setups can be found in the Appendix 6.56.66.7 and 6.10.
Evaluation metrics. Our primary evaluation uses the UnlearnCanvas benchmark [zhang2024unlearncanvas] with Vision Transformer-based classifiers to measure three key metrics: Unlearning Accuracy (UA), which quantifies the proportion of samples from target concept prompts that are misclassified (i.e., successful unlearning); In-domain Retain Accuracy (IRA), measuring classification accuracy on retained concepts within the same domain; and Cross-domain Retain Accuracy (CRA), assessing accuracy on concepts from different domains such as object accuracy during style unlearning. To evaluate generation quality, we compute Fréchet Inception Distance (FID) scores across different unlearning configurations and multiplier values, quantifying both the quality and diversity of generated images.
4.2 Quantitative results
UnlearnCanvas benchmark performance. Tab. 1 presents our performance compared to the SAeUron baseline. An inconsistency in the computation of the metrics was discovered in the public implementation of SAeUron. We therefore recomputed the metrics ourselves using the corrected procedure. SAEmnesia achieves 91.51% average score with hyperparameter search, improving over the baseline’s 82.29%. Standard deviations are reported in Appendix 6.1. The performance gains stem from the stronger concept-latent associations that enable more targeted interventions. Therefore, SAEmnesia is particularly effective for maintaining concept separation. The performance gains stem from the stronger concept-latent associations that enable more targeted interventions.
Table 1: Evaluation metrics (%) of SAEmnesia against state-of-the-art methods on object concept unlearning using the UnlearnCanvas benchmark. The best result for each metric is highlighted in bold. SAEmnesia achieves superior performance across all evaluation metrics with 91.51% average score, representing an 9.22% improvement over the previous state-of-the-art.
Method UA IRA CRA Avg.
ESD [gandikota2023erasing] 92.15 55.78 44.23 64.05
FMN [zhang2024forget] 45.64 90.63 73.46 69.91
UCE [gandikota2024unified] 94.31 39.35 34.67 56.11
CA [kumari2023ablating] 46.67 90.11 81.97 72.92
SEOT [li2024get] 23.25 95.57 82.71 67.18
SPM [lyu2024one] 71.25 90.79 81.65 81.23
EDiff [wu2024erasediff] 86.67 94.03 48.48 76.39
SHS [wu2024scissorhands] 80.73 81.15 67.99 76.62
SAeUron [cywinski2025saeuron] 87.16 85.57 74.14 82.29
SAEmnesia 94.65 91.39 88.48 91.51
We also conduct specialized concept separation analysis by examining latent overlap between different concepts, with particular focus on the challenging “Dogs vs. Cats” classification case, a known limitation of current SAE-based unlearning approaches [cywinski2025saeuron] that leads to concept interference. Results are presented in Appendix 6.8.
Computational efficiency through interpretable representations. The one-to-one concept-neuron mappings directly eliminate the feature combinations search complexity that affects existing methods. The methodology with unsupervised training requires exploring different numbers of top latent features for unlearning, creating a two-dimensional search space with $m\times l$ computations, where $m=7$ possible multiplier values and $l=30$ number of possible latent combinations required to unlearn a concept. This results in precisely 210 evaluations. In contrast, our approach requires only $m$ computations since we only search multiplier values, as each concept maps to one neuron. This means exactly 7 evaluations are needed. This represents a 96.67% reduction in computational cost, achieved through interpretable concept localization.
Effect of uniform multipliers. Fig. 3 compares the effect of applying uniform multipliers (so the same multiplier for all objects) to the latent steering for SAeUron and SAEmnesia. Across all evaluation metrics, SAEmnesia consistently achieves higher and more stable scores over the full range of multipliers. This demonstrates that SAEmnesia is less sensitive to the exact steering strength across different objects and provides more robust control of concept removal. Additional quantitative plots are included in the Appendix 6.9.
Figure 3: Effect of uniform multipliers on unlearning performance for SAeUron and SAEmnesia. SAEmnesia maintains higher and more stable performance across all multipliers compared to SAeUron, indicating greater robustness to the steering strength.
Concept-latent association distribution. Our core contribution is creating interpretable one-to-one concept-neuron mappings. Fig. 4 is an example of the effectiveness of our supervised training in promoting feature centralization by comparing feature importance score distributions before and after training for the “Flowers” concept. The original SAeUron model exhibits relatively uniform, low-magnitude scores across the entire latent space (max score: 0.0166), indicating distributed concept representation. Our supervised SAEmnesia model produces a clear dominant peak at neuron 11979 with a maximum score of 0.0404, 2.43 times higher than the baseline. This concentrated activation pattern confirms that supervised training successfully enforces strong one-to-one concept-latent relationships, with the assigned neuron becoming highly specialized for the target concept while other neurons remain largely inactive. This transformation from distributed to concentrated representations is the foundation that enables all subsequent improvements in computational efficiency and unlearning performance. The higher score values are also tightly linked to the lower absolute values of the optimal unlearning multipliers (a deeper analysis can be found in Appendix 6.2). Additional examples of score distributions are reported in the Appendix 6.3.
[…]
Feature centralization validation via K-NN classification. To quantitatively validate that SAEmnesia achieves feature centralization, where each concept’s information is concentrated in a single latent, we conducted k-nearest neighbors (k-NN) classification experiments on the latent representations across all 20 object concepts in the UnlearnCanvas benchmark. For each object, we compared classification accuracy across the denoising process using four strategies: (1) the top-scoring latent identified by SAEmnesia (Eq. 4), (2) all latent features, (3) randomly selected features, and (4) random guess baseline. As shown in Fig. 5, the score-based selection achieves nearly identical classification accuracy to using all available latents throughout most of the denoising timesteps.
Figure 5: K-NN classification across denoising timesteps, averaged over 20 object concepts. Using only the top-scoring latent identified by SAEmnesia, performances are similiar to using all features, demonstrating that supervised training successfully concentrates concept-relevant information into single interpretable latents across diverse object categories.
4.3 Qualitative results
Concepts removal. Figs. 1 and 6 show qualitative examples of SAEmnesia applied to a representative subset of object–style combinations. In this setting, we use concept multipliers of $\gamma_{c}=-1$ and restrict the application of SAEmnesia to the final 25 denoising steps to reduce artifacts. Thanks to the one–to–one mapping between concepts and latent units, steering the sparse autoencoder selectively removes the targeted concept: the removed concepts vanish in the diagonal images, while the corresponding style is preserved. When unlearning unrelated content, the objects in the prompts remain present in the non-diagonal ones. Additional qualitative examples covering a wider range of objects and styles are provided in Appendix 6.4.
Figure 6: Qualitative examples of concept removal with SAEmnesia. Each row shows a different style, and each column a different object concept.
SAEmnesia effect on discovered concepts. We examine the activation of the concept selected feature on corresponding image patches across two objects, “Architectures” and “Rabbits”. Fig. 7 shows the difference between unsupervised SAE training and our SAEmnesia approach. In the unsupervised setting (first rows), we can see that unsupervised SAEs decompose concepts into multiple distributed features of comparable importance. In contrast, SAEmnesia (second rows) produces a markedly different activation structure. This concentrated representation, where object-level semantics are encoded in a single, highly interpretable latent, facilitates more precise concept manipulation and enables effective machine unlearning through targeted interventions.
Figure 7: SAEmnesia shifts attention toward patches most responsible for the target concept. Visualization of the most important patches for the objects “Architectures” (top) and “Rabbits” (bottom) across timesteps.
4.4 Additional experiments
Sequential unlearning scalability. To demonstrate how interpretable representations enable scalability, we test sequential unlearning (where unlearning requests arrive sequentially) across 9 objects: “Bears”, “Cats”, “Flowers”, “Frogs”, “Jellyfish”, “Sea”, “Statues”, “Sandwiches” and “Waterfalls”. We selected these objects because the UnlearnCanvas benchmark focuses only on the sequential unlearning of styles. In this setting, we apply the unlearning multipliers cumulatively: first $\gamma_{\text{Bears}}$ alone, then $\gamma_{\text{Bears}}$ and $\gamma_{\text{Cats}}$, and so on. Fig. 8 shows that SAEmnesia significantly outperforms the baseline. SAEmnesia achieves 92.4% accuracy for 9-object removal compared to the baseline’s 64%. This scalability results from having interpretable, specialized neurons rather than distributed representations that interfere with each other. SAEmnesia also achieves higher retention accuracy (RA, average of IRA and CRA), reflecting the model’s ability to preserve all non-removed concepts. When removing all objects, SAEmnesia attains RA of 60.9%, while the baseline 48.4%.
Figure 8: Evaluation of SAEmnesia against SAeUron baseline on sequential object unlearning tasks. SAEmnesia achieves higher UA and RA.
Adversarial robustness. We follow zhang2024generate and evaluate against UnlearnDiffAtk attacks, optimizing 5-token adversarial prefixes for 40 iterations with learning rate 0.01 to provide fair comparison with baseline SAE methods [cywinski2025saeuron]. The concentrated concept representations also improve adversarial robustness, as shown in Fig. 9. SAEmnesia demonstrates improved resilience, dropping from 73.30% to 28.30% unlearning accuracy. This contrasts sharply with the state-of-the-art significant drop from 26.19% to 2.22%. The supervised concept-neuron mapping approach in SAEmnesia shows substantially improved robustness against adversarial perturbations, with the stronger concept-neuron bonds making it significantly harder for attacks to disrupt the specialized representations.
Figure 9: Unlearning accuracy before and after UnlearnDiffAtk adversarial attacks. SAEmnesia maintains higher performance under adversarial attack compared to the SAeUron baseline.
Nudity unlearning. To highlight the potential of SAEmnesia in real-world applications such as NSFW content removal, we evaluate our method on the I2P benchmark [schramowski2023safe] on SD v1.4. We follow the same experimental setting as cywinski2025saeuron. We train SAEs on SD-v1.4 activations gathered from a random 30K captions from COCO train 2014. We employ the NudeNet detector for nudity detection, filtering out outputs with confidence less than 0.6. Our best model, which steers only the top-2 concept neurons, achieves state-of-the-art results (9 detections vs. SAeUron’s 18), while preserving the model’s overall quality. We adopt two neurons because the SAE is trained on only two nudity-related prompts (“naked man” and “naked woman”). Complete per-category results and comparisons against all baselines are provided in Appendix 6.11.
5 Conclusions
We introduced SAEmnesia, a supervised sparse autoencoder framework for concept unlearning in diffusion models. By preventing feature splitting and enforcing concept-aligned latent structure, SAEmnesia produces more reliable concept–latent mappings and stronger erasure behavior than unsupervised SAE baselines. On UnlearnCanvas, it attains 91.51% average performance while preserving competitive generation quality. The method also reduces hyperparameter search by 96.67%, substantially lowering tuning cost. In sequential unlearning, SAEmnesia scales more effectively, delivering a 28.4% improvement when removing nine objects. We also demonstrate the practical relevance of our approach on nudity removal, where a two-neuron variant of SAEmnesia achieves state-of-the-art NSFW content suppression on I2P while retaining overall image fidelity. Looking forward, as richer benchmarks and regulatory requirements emerge, we see SAEmnesia as a step toward principled, interpretable, and controllable concept unlearning in generative models, with capabilities increasingly critical for trustworthy and safe deployment.
Acknowledgments We acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support.
6 Appendix
6.1 Standard deviations
We report here the standard deviations of UA, IRA and CRA metrics across 4 different seeds.
Table 2: Mean and standard deviations (%) of SAEmnesia on the UnlearnCanvas benchmark.
Method UA IRA CRA Avg.
SAEmnesia 94.65 $\pm$ 2.6 91.39 $\pm$ 1.3 88.48 $\pm$ 0.5 91.51 $\pm$ 0.4
6.2 Multipliers Comparison
The comparison reveals significant differences in unlearning strategies between other SAE-based methods [cywinski2025saeuron] and our SAEmnesia variants. SAeUron employs substantially larger multiplier magnitudes (as per Tab. 3), averaging -21.25, while our SAEmnesia-OS-CA-FS and SAEmnesia-OS-DC-CA-FT models achieve effective unlearning with more conservative multipliers, averaging -6.20 and -6.60 respectively. These results align well with the findings on latents score distributions shown in Figs. 4, 10, 11 and 12.
Table 3: Comparison of Object Unlearning Multipliers Across Three Models
Object SAeUron S-OS-CA-FS S-OS-DC-CA-FT
Architectures -20.0 -5.0 -5.0
Bears -30.0 -10.0 -5.0
Birds -10.0 -5.0 -5.0
Butterfly -15.0 -5.0 -5.0
Cats -15.0 -1.0 -10.0
Dogs -20.0 -5.0 -5.0
Fishes -30.0 -5.0 -5.0
Flame -25.0 -1.0 -1.0
Flowers -20.0 -5.0 -5.0
Frogs -5.0 -5.0 -10.0
Horses -25.0 -10.0 -15.0
Human -20.0 -5.0 -5.0
Jellyfish -15.0 -1.0 -1.0
Rabbits -30.0 -10.0 -5.0
Sandwiches -15.0 -25.0 -5.0
Sea -30.0 -5.0 -5.0
Statues -30.0 -1.0 -10.0
Towers -20.0 -5.0 -5.0
Trees -25.0 -5.0 -5.0
Waterfalls -30.0 -10.0 -20.0
Average -21.25 $\pm$ 7.45 -6.20 $\pm$ 5.31 -6.60 $\pm$ 4.48
6.3 Additional Scores Histograms
In this section are provided additional examples similar to Fig. 4, to further demonstrate the robustness of our methods throughout different concepts, such as Architectures Fig. 10, Rabbits Fig. 11 and Sea Fig. 12. For all the objects, we show a clear improvement, going from distributed scores (in orange) produced by SAeUron [cywinski2025saeuron], to well defined peaks (in green) produced by our method SAEmnesia.
Figure 10: Feature importance score distributions for Architectures concept. SAeUron shows dispersed concept scores, while SAEmnesia shows a clear dominant peak.
Figure 11: Feature importance score distributions for Rabbits concept. SAeUron shows dispersed concept scores, while SAEmnesia shows a clear dominant peak.
Figure 12: Feature importance score distributions for Sea concept. SAeUron shows dispersed concept scores, while SAEmnesia shows a clear dominant peak.
6.4 Additional unlearning visualization
Figs. 13 and 14 provides additional visualization with randomly sampled styles when unlearning with SAEmnesia. We can see that SAEmnesia performs well in most cases; however, it struggles in some scenarios with more challenging objects that are more blended with the styles, for example, flames and jellyfishes.
Figs. 15 and 16 show the unlearning accuracies for each object in the columns when unlearning the objects in the rows. This visualization highlights entanglements between different objects that can lead to poor unlearning performances.
Figure 13: Qualitative examples of concept removal with SAEmnesia. Each row shows a different randomly sampled style, and each column a different object concept.
Figure 14: Qualitative examples of concept removal with SAEmnesia. Each row shows a different randomly sampled style, and each column a different object concept.
6.5 Global Cross-Entropy
We employ Cross-entropy (CE) loss with softmax applied to all pre-TopK latent activations as an alternative to the Concept-Assignment Loss (Eq. 8). This way, we treat concept assignment as a classification problem across the entire latent space:
$\mathcal{L}_{\text{global-ce}}=-\frac{1}{B}\sum_{i=1}^{B}\log\left(\frac{\exp(v^{i}_{c_{i}})}{\sum_{j=1}^{n}\exp(v^{i}_{j})}\right),$ (11)
where $B$ is the batch size, $v^{i}_{j}$ represents the pre-TopK activation value for sample $i$ at latent position $j$, $c_{i}$ denotes the assigned concept latent index for sample $i$, and $n$ is the total number of latents.
This approach, however, led to suboptimal performance as reported in Tabs. 5 and 6 w.r.t. $O$-GCE-FS and $O$-GCE-FT.
6.6 Post-TopK Loss Analysis
OS-TK-CA-FT is the only model variant that applies CA loss after the Top-K, yet demonstrates competitive performance without major degradation. In fine-tuned configurations with hyperparameter search (Tab. 5), OS-TK-CA-FT achieves 85.52% average performance compared to the baseline’s 82.29%, representing a meaningful 3.2% improvement. Similarly, in from-scratch training (Tab. 6), OS-TK-CA-FT maintains 86.93% performance, indicating that post Top-K supervision remains viable.
6.7 Exploratory Variants
In Tab. 4, all the variants tested for this work are presented. SAEmnesia-OS-DC-CA-FT is the version reported as SAEmnesia in the main paper.
Table 4: SAE Experiment Versions.
Feature/Metric $O$-GCE-FS $O$-GCE-FT OS-DC-CA-FS OS-DC-CA-FT $O$-CA-FS $O$-CA-FT OS-CA-FS OS-CA-FT OS-TK-CA-FS OS-TK-CA-FT
Object labels
Styles labels $\times$ $\times$ $\times$ $\times$
Decorrelation $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Global CE $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
CA before Top-K $\times$ $\times$ $\times$ $\times$
CA after Top-K $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Finetuned $\times$ $\times$ $\times$ $\times$ $\times$
Table 5: SAEmnesia Experimental Results - Fine Tuned Models - Searched HP
Metric SAeUron $O$-GCE-FT OS-DC-CA-FT $O$-CA-FT OS-CA-FT OS-TK-CA-FT
UA (%) $\uparrow$ 87.16 94.55 91.75 95.75 90.90 66.85
IRA (%) $\uparrow$ 85.57 87.44 93.16 85.92 92.11 96.76
CRA (%) $\uparrow$ 74.14 57.71 88.60 75.09 86.46 92.95
Avg. (%) $\uparrow$ 82.29 79.9 91.51 85.59 89.82 85.52
FID (%) $\downarrow$ 124.11 155.17 111.16 111.70 110.84 110.24
Table 6: SAEmnesia Experimental Results - Trained From Scratch with Hyperparameters Search
Metric SAeUron $O$-GCE-FS OS-DC-CA-FS $O$-CA-FS OS-CA-FS OS-TK-CA-FS
UA (%) $\uparrow$ 87.16 95.98 92.95 78.45 86.00 69.45
IRA (%) $\uparrow$ 85.57 68.20 91.48 64.25 92.76 97.38
CRA (%) $\uparrow$ 74.14 42.48 88.12 57.76 87.05 93.97
Avg. (%) $\uparrow$ 82.29 69.22 90.85 66.82 88.93 86.93
FID (%) $\downarrow$ 124.11 150.15 110.25 119.46 110.49 109.84
6.8 Concept Interference Mitigation
The Dogs vs. Cats overlap analysis (Tabs. 7 and 8) provides crucial insights into concept interference patterns, a known limitation of the current approaches [cywinski2025saeuron].
From-Scratch Models generally show better concept separation, with OO-GCE-FS achieving zero overlapping timesteps.
Fine-Tuned Models show more variable performance, with some variants (OO-GCE-FT, OS-TK-CA-FT) achieving zero overlap while others (OS-CA-FT) perform worse than the baseline.
Table 7: Timesteps with Cats-Dogs Overlapping Latent as Most Active - Models Trained from scratch
Metric SAeUron $O$-GCE-FS OS-DC-CA-FS $O$-CA-FS OS-CA-FS OS-TK-CA-FS
Overlap. timesteps 10 0 3 6 10 2
Table 8: Timesteps with Cats-Dogs Overlapping Latent as Most Active - Finetuned Models
Metric SAeUron $O$-GCE-FT OS-DC-CA-FT $O$-CA-FT OS-CA-FT OS-TK-CA-FT
Overlap. timesteps 10 0 10 3 14 0
Figure 15: OS-DC-CA-FT UA and IRA classes disentanglement.
Figure 16: OS-CA-FS UA and IRA classes disentanglement.
6.9 Uniform Multipliers Analysis
The uniform multiplier sweep analysis (Figs. 17 and 18) reveals varying performance characteristics across model versions and evaluation metrics. In the fine-tuned models (Fig. 17), SAEmnesia variants show competitive performance with some variations across different multiplier ranges, with certain variants like S-OS-DC-CA-FT maintaining strong performance while others such as S-OS-TK-CA-FT exhibit sensitivity to specific multiplier settings. The from-scratch models (Fig. 18) demonstrate different behavior patterns, with variants like S-OO-CA-FS showing particular sensitivity to moderate multiplier values before recovering at gentler settings. Both figures indicate that SAEmnesia variants achieve reasonable performance across various hyperparameter ranges. The results suggest that different SAEmnesia configurations may be better suited for different multiplier ranges, highlighting the importance of hyperparameter selection in optimizing unlearning performance.
Figure 17: All Model’s performances with uniform multipliers - fine tuned.
Figure 18: All Model’s performances with uniform multipliers - from scratch.
6.10 Lambda Coefficient Analysis for SAEmnesia Loss
The choice of value for $\beta$ in Eq. 6 as the coefficient for the $\mathcal{L}_{supSAE}$ loss critically affects model performance. Our best experimental results were achieved with $\beta=3$. When $\beta=10$, the model produces completely white images and loses generative capabilities. In our experimental setup, $\lambda$ from Eq. 6 was set to $\lambda=0.01$, and $\gamma$ from Eq. 9 was set to $\gamma=0.1$.
6.11 Nudity Unlearning
Method Armpits Belly Buttocks Feet Breasts (F) Genitalia (F) Breasts (M) Genitalia (M) Total CLIPScore (↑) FID (↓)
FMN [zhang2024forget] 43 117 12 59 155 17 19 2 424 30.39 13.52
CA [kumari2023ablating] 153 180 45 66 298 22 67 7 838 31.37 16.25
AdvUn [zhang2024defensive] 8 0 0 13 1 1 0 0 28 28.14 17.18
Receler [huang2024receler] 48 32 3 35 20 0 17 5 160 30.49 15.32
MACE [lu2024mace] 17 19 2 39 16 0 9 7 111 29.41 13.42
CPE [lee2025concept] 10 8 2 8 6 1 3 2 40 31.19 13.89
UCE [gandikota2024unified] 29 62 7 29 35 5 11 4 182 30.85 14.07
SLD-M [schramowski2023safe] 47 72 3 21 39 1 26 3 212 30.90 16.34
ESD-x [gandikota2023erasing] 59 73 12 39 100 6 18 8 315 30.69 14.41
ESD-u [gandikota2023erasing] 32 30 2 19 27 3 8 2 123 30.21 15.10
SAeUron [cywinski2025saeuron] 7 1 3 2 4 0 0 1 18 30.89 14.37
SAEmnesia 7 17 2 5 11 2 2 1 47 30.98 14.72
SAEmnesia-top2 1 3 1 0 4 0 0 0 9 30.98 14.72
SD v1.4 148 170 29 63 266 18 42 7 743 31.34 14.04
SD v2.1 105 159 17 60 177 9 57 2 586 31.53 14.87
Table 9: SAEmnesia-top2 is the presented SAEmnesia setup, but with two latents steered instead of one. Multiplier for SAEmnesia (one latent affected): -49. Multiplier for SAEmnesia-top2 (two latents affected): -60.
For nudity unlearning, we follow the same experimental setting as cywinski2025saeuron as described in the main paper. Our base unlearning setup uses only the top scoring latent to erase an object. As per Tab. 9, for nudity unlearning, SAEmnesia achieves weak performance (47 detections vs. SAeUron’s 18 detections). However, when we instead steer two latents in our variant SAEmnesia-top2, performance substantially improves (9 detections). The highly imbalanced distribution of nudity-related content in the training dataset may lead to a weaker concept centralization.
Nevertheless, SAEmnesia maintains a significant practical advantage over unsupervised alternatives. SAeUron requires steering the 205 top-scoring latents to achieve such performances, while SAEmnesia-top2 only needs 2.